# Value Gradients with Action Adaptive Search Trees in Continuous (PO)MDPs

### v1

@@ B.3 Evaluations
$n_{D}^{sims,POMCPOW}=500\cdot$0.08$\cdot J_{D},$ (110)
which was empirically found to produce similar runtimes to PFT-DPW in the chosen domain.
In order to save computation time and not compute a partially observable rollout, the rollout computation for a new particle belief was done by drawing $s_{0}$ and 10 additional particles from the belief. The rollout actions were computed based on $s_{0}$, but the return was the mean of returns from the 10 particles. This mechanism allows to approximate a partially observable rollout without receiving any future observations. For POMCPOW, since every new node is expanded initially only with a single particle, the rollout was computed with that particle. This is a disadvantage for POMCPOW as it made the rollout returns less meaningful, however it seemd to have achieved better results than PFT-DPW despite that.
The algorithm hyperparameters were first determined by cross entropy optimization (CE) [?], largely following [?], using Gaussian distributions over the continuous DPW hyperparameters:
  1. 1.
    UCT exploration bonus $c$.
  2. 2.
    Action progressive widening parameters $k_{a}$ and $\alpha_{a}$.
  3. 3.
    Observation progressive widening parameters $k_{o}$ and $\alpha_{o}$.
For AGMCTS, we also optimized over the Adam learning rate $\eta_{Adam}$, and the action update distance threshold $T_{d_{a}}$. We searched for hyperparameters for each dimension while using rollout policy noise $\sigma_{r}=0.1$, and the results for $\sigma_{r}=0.2,0.3$ were obtained using the same hyperparameters of the same dimension. We used 150 parameter samples, 40 simulations for each parameter sample to determine its mean return, used 30 elite samples to fit the new distribution, with smoothing the new distribution’s parameters:
where we chose $\alpha_{\mu}=0.8,\alpha_{\Sigma}=0.5$.
We performed 50 CE iterations for each algorithm in each dimension. The starting mean for $D=2$ was the same for all algorithms, and for $D=3,4$ we used the mean of the best hyperparameters found for $D-1$. The hyperparameters chosen were those from the CE iteration achieving the maximal mean of mean returns over the 30 elite samples. The parameters used for each algorithm are summarized in Table 3. After choosing the parameters, we reported the mean return over 1000 scenarios, with the standard error of the mean. Table 4 summarizes the results with the individual mean runtimes for each algorithm in each scenario.
D $\sigma_{r}$ POMCPOW time (s) PFT-DPW time (s) AGMCTS time (s)
2 0.1 $\boldsymbol{6.62}\pm 0.07$ 0.114 $5.71\pm 0.08$ 0.126 $\boldsymbol{6.67}\pm 0.06$ 1.665
2 0.2 $\boldsymbol{6.23}\pm 0.07$ 0.122 $4.74,0.08$ 0.129 $5.96\pm 0.08$ 1.730
2 0.3 $\boldsymbol{5.61}\pm 0.08$ 0.129 $4.01\pm 0.09$ 0.129 $\boldsymbol{5.49}\pm 0.09$ 1.838
3 0.1 $3.42\pm 0.08$ 0.216 $3.63\pm 0.08$ 0.299 $\boldsymbol{5.47}\pm 0.07$ 2.442
3 0.2 $1.87\pm 0.08$ 0.240 $1.79\pm 0.08$ 0.296 $\boldsymbol{3.82}\pm 0.1$ 2.416
3 0.3 $1.06\pm 0.07$ 0.253 $0.29\pm 0.06$ 0.309 $\boldsymbol{1.58}\pm 0.1$ 2.389
4 0.1 $2.58\pm 0.07$ 0.548 $2.36\pm 0.07$ 0.683 $\boldsymbol{4.26},0.08$ 10.576
4 0.2 $0.53\pm 0.05$ 0.640 $0.2\pm 0.05$ 0.707 $\boldsymbol{1.53}\pm 0.09$ 8.303
4 0.3 $\boldsymbol{-0.36}\pm 0.03$ 0.660 $-0.86\pm 0.02$ 0.736 $-1.13\pm 0.04$ 5.778
Table 4: Performance results for the different algorithms at D-Continuous Light Dark for $D=2,3,4$ dimensions. AGMCTS and PFT-DPW were run for 500 simulations with particle counts given by (108), and POMCPOW for an approximately equivalent time to PFT-DPW according to (110). The rollout policy’s standard deviation was $\sigma_{r}=0.1$ during hyperparameter tuning, and then the algorithms were also tested for $\sigma_{r}=0.2,0.3$. The reported results are mean returns over 1000 scenarios $\pm$ standard error. AGMCTS achieved the best performance in most scenarios.

### v3

@@ C.2.2 Evaluation
$n_{pomcpow,D}^{sims}=500\cdot$0.08$\cdot J_{D}.$ (159)
Additional parameters that were manually set in all dimensions were:
[…]
$T_{d_{a}}^{\max}=\infty,\quad K_{\text{child}}^{\min}=1,\quad K_{\text{child}}^{visits}=1,\quad K_{b}^{\nabla}=5.$ (160)
In this domain, exponential step size decay was used.
Table 3: Performance and mean time results at 2D-Continuous Light-Dark.
$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
1024 $3.60\pm 0.11$ 0.018 50 $3.94\pm 0.09$ 0.014 $\boldsymbol{4.27}\pm 0.09$ 0.018
1823 $4.63\pm 0.10$ 0.016 89 $4.62\pm 0.09$ 0.022 $\boldsymbol{5.28}\pm 0.09$ 0.041
3236 $5.21\pm 0.10$ 0.033 158 $5.06\pm 0.09$ 0.041 $\boldsymbol{6.06}\pm 0.08$ 0.201
5755 $\boldsymbol{6.28}\pm 0.08$ 0.058 281 $5.52\pm 0.08$ 0.073 $\boldsymbol{6.26}\pm 0.07$ 0.593
10240 $\boldsymbol{6.62}\pm 0.07$ 0.114 500 $5.71\pm 0.08$ 0.126 $\boldsymbol{6.67}\pm 0.06$ 1.665
Table 4: Performance and mean time results at 3D-Continuous Light-Dark.
$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
2048 $\boldsymbol{2.37}\pm 0.09$ 0.025 50 $\boldsymbol{2.37}\pm 0.08$ 0.031 $2.01\pm 0.08$ 0.328
3645 $2.36\pm 0.09$ 0.036 89 $\boldsymbol{2.78}\pm 0.08$ 0.052 $\boldsymbol{2.81}\pm 0.10$ 0.162
6472 $\boldsymbol{3.39}\pm 0.09$ 0.065 158 $3.12\pm 0.08$ 0.089 $3.09\pm 0.10$ 0.432
11510 $3.60\pm 0.08$ 0.116 281 $3.44\pm 0.08$ 0.164 $\boldsymbol{5.46}\pm 0.07$ 1.064
20480 $3.42\pm 0.08$ 0.216 500 $3.63\pm 0.08$ 0.299 $\boldsymbol{5.47}\pm 0.07$ 2.442
Table 5: Performance and mean time results at 4D-Continuous Light-Dark.
$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
4096 $\boldsymbol{1.48}\pm 0.07$ 0.052 50 $0.85\pm 0.07$ 0.065 $0.65\pm 0.06$ 0.132
7291 $\boldsymbol{2.13}\pm 0.07$ 0.089 89 $1.47\pm 0.07$ 0.116 $1.34\pm 0.09$ 0.318
12943 $\boldsymbol{2.15}\pm 0.08$ 0.157 158 $1.98\pm 0.07$ 0.213 $1.79\pm 0.09$ 1.031
23020 $2.44\pm 0.07$ 0.286 281 $2.13\pm 0.07$ 0.382 $\boldsymbol{3.75}\pm 0.09$ 3.024
40960 $2.58\pm 0.07$ 0.548 500 $2.36\pm 0.07$ 0.683 $\boldsymbol{4.26}\pm 0.08$ 10.57
C.3 Mountain Car and Hill Car
C.3.1 Problem Description
A car is placed randomly in a valley between two hills, and the goal is to reach the top of the right hill. The car’s action is its acceleration left or right, and must gain momentum in order to reach the goal. In this domain, the state $s=(x,v)\in\mathbb{R}^{2}$ is the car’s position and velocity, $\mathcal{A}=[-a_{\max},a_{\max}]\subset\mathbb{R}$ is the action space.
In the POMDP version, $z\in\mathbb{R}$ is the observation of the car’s position with noise. It is given by $z\sim\mathcal{N}(x,\sigma_{O}^{2})$, where $\sigma_{O}=0.03$.
The car’s position is bounded to $[x_{\min},x_{\max}]$, and the velocity is bounded to $[-v_{\max},v_{\max}]$.
If the car reaches the goal $x\geq x_{\max}$, it receives a reward of $100$. If $x<x_{\min}$, or $\left|{v}\right|\geq v_{\max}$, a penalty reward of $-100$ is attained and the scenario terminates. Otherwise, the car obtains a reward of $-0.1$ at each time step that does not terminate. The reward function is a function of the posterior state, i.e. $r(s,a,s^{\prime})=r(s^{\prime})$. In Mountain Car, the scenario terminates after $L=200$ time steps, while in Hill Car it terminates after $L=30$ time steps. The discount in both domains is $\gamma=0.99$.
The rollout policy is calculated by choosing $a_{\max}$ when $v>0$ and $-a_{\max}$ otherwise.
In both domains, the generative transition model is given by an input action noise to the deterministic dynamics:
$\nu\sim\mathcal{N}(0,\sigma_{\nu}^{2}),$ (161)
$\tilde{a}=\text{clip}(a+w,-a_{\max},a_{\max}),$ (162)
$s^{\prime}=f(s,\tilde{a},s^{\prime}),$ (163)
where $\sigma_{\nu}=0.1$ for both domains.
When calculating the transition density in this domain, we do it based on the input noise to a deterministic simulator case via the Area Formula. For each $(s,a^{\prime},s^{\prime})$ tuple we calculate the generating action noise $\nu$, depending on the scenario. If the obtained $\nu$ is outside of the bounds (beyond numerical errors) $[L,R]$ where $L=-a_{\max}-a$ and $R=a_{\max}-a$, then the density is 00. Otherwise, because of the clip operation, we calculate its density by the following mixture:
$p(\nu\mid s,a)=\begin{cases}CDF(\mathcal{N}(0,\sigma_{\nu}^{2}),L),&\text{if }\nu\leq L\\ 1-CDF(\mathcal{N}(0,\sigma_{\nu}^{2}),R),&\text{if }\nu\geq R\\ PDF(\mathcal{N}(0,\sigma_{\nu}^{2}),\nu),&\text{otherwise}.\end{cases}$ (164)
Afterwards, we calculate the transition model density by the Area Formula according to the case of input noise to a deterministic simulator.
C.3.2 Mountain Car Transition Model
In this domain, we set $x_{\min}=-1.5$, $x_{\max}=0.5$, $v_{\max}=0.05$, $a_{\max}=1.0$. The deterministic dynamics are calculated by:
$v^{\prime}=v+0.001\cdot\tilde{a}-0.0025\cos(3x),$ (165)
$x^{\prime}=x+v^{\prime}.$ (166)
In order to calculate the transition density, for given $s,a^{\prime},s^{\prime}$ we invert the dynamics to obtain $\nu$.
C.3.3 Hill Car Transition Model
In this domain, we set $x_{\min}=-1.0$, $x_{\max}=1.0$, $v_{\max}=2.5$, $a_{\max}=4.0$. The deterministic dynamics are calculated by integrating the car’s continuous-time dynamics with an ODE solver over the time period $t_{0}$ to $t_{1}$. To limit the computational complexity, we chose crudely $\delta t=0.01$ for the integration, while $t_{1}-t_{0}=0.1$ was the time step between consecutive time steps of the MDP/POMDP scenario. The integration method chosen was Tsitouras 5/4 Runge-Kutta method.
The continuous-time dynamics are given by:
$\dot{v}=\left(\frac{a}{m}-g\cdot h^{\prime}(x)-v^{2}\cdot h^{\prime}(x)\cdot h^{\prime\prime}(x)\right)\frac{1}{1+h^{\prime}(x)^{2}},$ (167)
$\dot{x}=v,$ (168)
where $m=1$, $g=9.81$, and the hill-curve function $h(x)$ is given by:
$h(x)=\begin{cases}x^{2}+x,&\text{if }x<0\\ \frac{x}{\sqrt{1.0+5.0*x^{2}}},&\text{otherwise}.\end{cases}$ (169)
Because the input noise is only based on the action, we calculate $\nu$ by caching simulator outputs $(s,a,\nu,s^{\prime})$ tuples for which $s^{\prime}=f(s,\tilde{a})$. For a new tuple $(s,a^{\prime},s^{\prime})$, we calculate $\nu^{\prime}=\tilde{a}-a^{\prime}$ for the cached $\tilde{a}$. The jacobian of the dynamics is calculated by ForwardDiff.jl and SciMLSensitivity.jl overloaded gradient function for the ODE solver.
C.3.4 Evaluation
DPW, PFT-DPW and AGMCTS were given a simulation budget of $n^{sims}=500$ simulations. In the POMDP settings, the particle count was set to $J=30$, and we used $J^{PF}=200$ for inference between planning steps. POMCPOW simulation budget was set by the formula:
$n_{pomcpow}^{sims}=500\cdot$0.08$\cdot J.$ (170)
In this domain, we used linearized weight updates for AGMCTS MIS action update. Exponential step size decay was not used.
Table 6: Hyperparameters used in evaluations of AGMCTS in the Mountain/Hill Car domains.
Parameter Mountain/MDP Mountain/POMDP Hill/MDP Hill/POMDP
$c$ 0.0 0.001 132.24 132.83
$k_{a}$ 6.876 4.558 6.552 8.657
$\alpha_{a}$ 0.619 0.698 0.532 0.490
$k_{o}$ 0.292 0.379 5.375 6.050
$\alpha_{o}$ 0.385 0.382 0.203 0.130
[…]
$T_{d_{a}}^{\min}$ 0.0 0.0 0.0 0.0
$J$ 1 30 1 30
$K_{rollout}$ 1 5 1 5
$K_{opt}$ 3 3 3 3
$T_{d_{a}}^{\max}$ 0.1 0.1 0.1 0.1
$T_{\omega}^{add}$ 1.0 0.99 1.0 0.99
[…]
$K_{\text{child}}^{\min}$ 2 2 2 2
$K_{\text{child}}^{visits}$ 1 1 1 1
$K_{b}^{\nabla}$ 3 - 3
$K_{O}^{\nabla}$ linearized linearized linearized linearized
$n^{sims}$ 500 500 500 500
Table 7: Hyperparameters used in evaluations of DPW/PFT-DPW in the Mountain/Hill Car domains.
Parameter Mountain/MDP Mountain/POMDP Hill/MDP Hill/POMDP
$c$ 92.148 146.08 132.24 119.28
$k_{a}$ 6.672 5.625 6.552 7.386
$\alpha_{a}$ 0.581 0.824 0.532 0.528
$k_{o}$ 0.277 1.049 5.375 1.256
$\alpha_{o}$ 0.454 0.415 0.203 0.588
$J$ 1 30 1 30
$K_{rollout}$ 1 5 1 5
$n^{sims}$ 500 500 500 500
Table 8: Hyperparameters used in evaluations of POMCPOW in the Mountain/Hill Car domains.
Parameter Mountain/POMDP Hill/POMDP
$c$ 59.585 90.960
$k_{a}$ 4.082 6.944
$\alpha_{a}$ 0.640 0.209
$k_{o}$ 0.520 5.052
$\alpha_{o}$ 0.197 0.845
$n^{sims}$ 2739 2739
Table 9: Performance and mean time results at Mountain Car MDP.
$n^{sims}$ DPW time(s) AGMCTS time(s)
50 $\boldsymbol{-52.00}\pm 0.17$ 0.001 $\boldsymbol{-52.00}\pm 0.17$ 0.0011
89 $\boldsymbol{-51.66}\pm 0.25$ 0.002 $\boldsymbol{-51.66}\pm 0.25$ 0.0030
158 $\boldsymbol{-48.00}\pm 0.58$ 0.004 $-51.05\pm 0.31$ 0.0053
281 $-45.43\pm 0.66$ 0.008 $\boldsymbol{29.43}\pm 0.07$ 0.0378
500 $24.94\pm 0.30$ 0.012 $\boldsymbol{29.52}\pm 0.066$ 0.0865
Table 10: Performance and mean time results at Mountain Car POMDP.
$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
274 $24.52\pm 0.33$ 0.007 50 $\boldsymbol{24.95}\pm 0.33$ 0.002 $\boldsymbol{25.29}\pm 0.32$ 0.003
487 $\boldsymbol{25.26}\pm 0.32$ 0.014 89 $24.45\pm 0.37$ 0.005 $24.11\pm 0.42$ 0.006
865 $\boldsymbol{25.78}\pm 0.28$ 0.025 158 $23.85\pm 0.41$ 0.009 $24.45\pm 0.41$ 0.011
1539 $25.00\pm 0.36$ 0.044 281 $23.29\pm 0.44$ 0.017 $\boldsymbol{26.46}\pm 0.24$ 0.022
2739 $24.24\pm 0.43$ 0.080 500 $23.16\pm 0.44$ 0.030 $\boldsymbol{26.65}\pm 0.19$ 0.040
Table 11: Performance and mean time results at Hill Car MDP.
$n^{sims}$ DPW time(s) AGMCTS time(s)
50 $\boldsymbol{-92.87}\pm 1.05$ 0.029 $\boldsymbol{-92.87}\pm 0.09$ 0.003
89 $\boldsymbol{-90.11}\pm 1.23$ 0.005 $\boldsymbol{-90.44}\pm 0.09$ 0.018
158 $-92.22\pm 1.10$ 0.009 $\boldsymbol{-90.60}\pm 0.08$ 0.016
281 $-89.30\pm 1.27$ 0.015 $\boldsymbol{17.25}\pm 0.07$ 0.060
500 $-88.66\pm 1.31$ 0.029 $\boldsymbol{51.82}\pm 0.06$ 0.260
Table 12: Performance and mean time results at Hill Car POMDP.
$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
274 $\boldsymbol{52.54}\pm 1.09$ 0.031 50 $-82.91\pm 1.56$ 0.030 $-77.49\pm 1.76$ 0.055
487 $\boldsymbol{53.98}\pm 1.00$ 0.054 89 $-75.58\pm 1.83$ 0.055 $-69.48\pm 1.99$ 0.063
865 $\boldsymbol{54.15}\pm 1.0$ 0.097 158 $-65.22\pm 2.09$ 0.098 $-67.16\pm 2.05$ 0.106
1539 $\boldsymbol{54.55}\pm 0.98$ 0.178 281 $-64.86\pm 2.10$ 0.183 $3.75\pm 2.42$ 0.227
2739 $\boldsymbol{55.65}\pm 0.90$ 0.311 500 $-36.53\pm 2.46$ 0.329 $51.79\pm 1.12$ 0.641
C.4 Lunar Lander
C.4.1 Problem Description
For this domain, we took the Julia implementation provided by Mern et al. [8].
The vehicle state is represented by a six dimensional tuple $(x,y,\theta,\dot{x},\dot{y},\omega)$, where $x$ and $y$ are the horizontal and vertical positions, $\theta$ is the orientation angle, $\dot{x}$ and $\dot{y}$ are the horizontal and vertical speeds, and $\omega$ is the angular rate. The action space $\mathcal{A}\subset\mathbb{R}^{3}$ is a three-dimensional continuous space defined by the tuple $(F_{x},T,\delta)$. $T$ is the main thrust which acts along the vehicle’s vertical axis through its center of mass, and is in the range $[0,15]$. $F_{x}$ is the corrective thrust, which acts along a horizontal axis offset from the center of mass by a distance $\delta$. $F_{x}$ is in the range $[-5,5]$ and $\delta\in[-1,1]$.
The transition model is computed as an additive Gaussian noise to the output of a deterministic transition function. The deterministic transition function is given by the following set of computations:
$f_{x}=\cos(\theta)\cdot F_{x}-\sin(\theta)\cdot T,$ (171)
$f_{z}=\cos(\theta)\cdot T+\sin(\theta)\cdot F_{x},$ (172)
$\tau=-\delta\cdot F_{x},$ (173)
$a_{x}=\frac{f_{x}}{m},\quad a_{z}=\frac{f_{z}}{m},\quad\dot{\omega}=\frac{\tau}{I},$ (174)
$v_{x}^{\prime}=v_{x}+a_{x}\cdot\Delta t+\epsilon_{1}\cdot Q_{4},$ (175)
$v_{z}^{\prime}=v_{z}+(a_{z}-9.0)\cdot\Delta t+\epsilon_{2}\cdot Q_{5},$ (176)
$\omega^{\prime}=\omega+\dot{\omega}\cdot\Delta t+\epsilon_{3}\cdot Q_{6},$ (177)
$x^{\prime}=x+v_{x}\cdot\Delta t,\quad z^{\prime}=z+v_{z}\cdot\Delta t,\quad\theta^{\prime}=\theta+\omega\cdot\Delta t,$ (178)
$s^{\prime}=\begin{bmatrix}x^{\prime}\\ z^{\prime}\\ \theta^{\prime}\\ v_{x}^{\prime}\\ v_{z}^{\prime}\\ \omega^{\prime}\end{bmatrix},$ (179)
where the noise variables $\epsilon_{i}$ are sampled from a standard normal distribution, $Q_{4}=0.1$, $Q_{5}=0.1$, $Q_{6}=0.01$. We took the values $m=1$, $I=10$, and $\delta t=0.4$. For computing the transition density, we used the Area Formula by solving the dynamics for the noise variables $\epsilon_{i}$ for given $(s,a^{\prime},s^{\prime})$ tuples.
In the POMDP variant, the vehicle makes noisy observations of its angular rate, horizontal speed, and above-ground height as measured by a distance sensor. The noise is Gaussian with a diagonal covariance matrix $\Sigma_{O}=\text{diag}(1.0^{2},0.01^{2},0.1^{2})$.
The reward function is defined as:
$r(s,a,s^{\prime})=\begin{cases}-1000,&\text{if }x\geq 15\vee\theta\geq 0.5\\ 100-x-\dot{y}^{2},&\text{if }y\leq 1\\ -1,&\text{otherwise}\end{cases}$ (180)
The initial vehicle state is sampled from a multivariate Gaussian with mean $\mu=(x=0,y=50,\theta=0,\dot{x}=0,\dot{y}=-10,\omega=0)$. The rollout policy was computed as a simple proportional rule: $a_{rollout}=(-0.1\cdot\dot{x},-0.1\cdot\dot{y},0)$.
C.4.2 Evaluation
DPW, PFT-DPW and AGMCTS were given a simulation budget of $n^{sims}=1000$ simulations. In the POMDP settings, the particle count was set to $J=150$, and we used $J^{PF}=2000$ for inference between planning steps. POMCPOW simulation budget was set by the formula:
$n_{pomcpow}^{sims}=1000\cdot$0.035$\cdot J.$ (181)
In this domain, we used linearized weight updates for AGMCTS MIS action update. Exponential step size decay was not used.
Specifically in the Lunar Lander MDP/POMDP domains, we observed that the planner was converging to a setting in which the gradients were practically never used. We tried repeating the optimization process via a constrained CE, in which rejection-sampling is used to constrain the parameter samples satisfy the following condition:
$\eta_{Adam}\cdot\frac{K_{opt}}{K_{\text{visits}}}\left(N-\text{ceil}\left(\left(\frac{K_{\text{child}}^{\min}}{k_{o}}\right)^{1/\alpha_{o}}\right)\right)\geq 0.01\cdot T_{d_{a}}^{\max}$ (182)
It seemed that forcing the planner to use the gradients was not beneficial in this domain, however in the unconstrained CE setting, the hyperparameters converged to a setting in which the gradients were not used, and its performance seemed to match that of PFT-DPW.
Table 13: Hyperparameters used in evaluations of AGMCTS in the Lunar Lander domains.
Parameter Lunar Lander MDP Lunar Lander POMDP
$c$ 61.55 112.14
$k_{a}$ 3.052 4.846
$\alpha_{a}$ 0.377 0.320
$k_{o}$ 0.114 1.347
$\alpha_{o}$ 0.047 0.273
[…]
$T_{d_{a}}^{\min}$ 0.0 0.0
$J$ 1 150
$K_{rollout}$ 1 5
$K_{opt}$ 1 5
$T_{d_{a}}^{\max}$ 0.1 0.1
$T_{\omega}^{add}$ 0.9 0.99
[…]
$K_{\text{child}}^{\min}$ 1 2
$K_{\text{child}}^{visits}$ 1 2
$K_{b}^{\nabla}$ 3
$K_{O}^{\nabla}$ linearized linearized
$n^{sims}$ 500 500
Table 14: Hyperparameters used in evaluations of DPW/PFT-DPW in the Lunar Lander domains.
Parameter Lunar Lander MDP Lunar Lander POMDP
$c$ 60.49 60.22
$k_{a}$ 1.421 2.687
$\alpha_{a}$ 0.595 0.436
$k_{o}$ 0.082 0.274
$\alpha_{o}$ 0.726 0.575
$J$ 1 150
$K_{rollout}$ 1 5
$n^{sims}$ 500 500
Table 15: Hyperparameters used in evaluations of POMCPOW in the Lunar Lander POMDP domain.
Parameter Lunar Lander POMDP
$c$ 66.04
$k_{a}$ 3.309
$\alpha_{a}$ 0.393
$k_{o}$ 0.354
$\alpha_{o}$ 0.433
$n^{sims}$ 12247
Table 16: Performance and mean time results at Lunar Lander MDP.
$n^{sims}$ DPW time(s) AGMCTS time(s)
100 $\boldsymbol{43.55}\pm 1.44$ 0.002 $\boldsymbol{44.15}\pm 1.06$ 0.023
178 $\boldsymbol{50.01}\pm 0.98$ 0.002 $\boldsymbol{48.49}\pm 1.03$ 0.010
316 $\boldsymbol{56.19}\pm 0.39$ 0.002 $53.88\pm 0.46$ 0.022
562 $\boldsymbol{58.29}\pm 1.02$ 0.004 $\boldsymbol{57.22}\pm 0.40$ 0.046
1000 $\boldsymbol{63.09}\pm 0.29$ 0.007 $60.02\pm 0.46$ 0.096
Table 17: Performance and mean time results at Lunar Lander POMDP.
$n^{sims}_{pomcpow}$ POMCPOW time(s) $n^{sims}$ PFT-DPW time(s) AGMCTS time(s)
1225 $\boldsymbol{50.96}\pm 3.02$ 0.015 100 $15.33\pm 5.09$ 0.0082 $-25.86\pm 6.38$ 0.026
2180 $\boldsymbol{53.06}\pm 2.88$ 0.023 178 $33.12\pm 3.95$ 0.0145 $0.872\pm 5.32$ 0.041
3870 $\boldsymbol{57.58}\pm 2.30$ 0.042 316 $40.44\pm 3.40$ 0.0262 $25.56\pm 4.12$ 0.078
6883 $\boldsymbol{58.48}\pm 2.28$ 0.075 562 $41.86\pm 3.55$ 0.0444 $37.68\pm 3.54$ 0.147
12247 $\boldsymbol{55.19}\pm 2.83$ 0.141 1000 $48.03\pm 2.95$ 0.0764 $42.98\pm 3.54$ 0.298